# Gold-medalist Performance in Solving Olympiad Geometry with AlphaGeometry2

### v2

@@ 7.2 Inference setup
A new problem is solved via the search algorithm described in section 6 with multiple search trees and multiple language models of different sizes. In contrast to AG1, we use top-k sampling with temperature $t=1.0$ and $k=32$. Note that a high temperature and multiple samples are essential for solving IMO problems. With the greedy decoding $t=0.0,k=1$, and no tree search, our models can solve only two problems out of 26 that require auxiliary constructions. Increasing the temperature to $t=1.0$ and using $k=32$ samples (without a search tree) allows our language models to solve 9 out of 26 problems. Lower temperatures $t<1.0$ do not produce diverse enough auxiliary constructions (see Figure 7), while higher temperatures result in the increasing number LM outputs with a wrong domain language syntax.
@@
Note that by definition, $S_{1}\subset S_{2}\subset S_{3}$. Once these three sets are computed, we serialize and concatenate them into a string called analysis string, using our domain specific language. This string is fed into the LM, together with the original problem statement as follows: <problem_statement> serialized($S_{1}$) serialized($S_{2}-S_{1}$) serialized($S_{3}-S_{2}$). In contrast, the input to the AG1 LM is simply <problem_statement>.
@@ 8 Results
Our main downstream metric is the solve rate on IMO geometry problems. There are a total of 45 geometry problems in 2000-2024 IMO, which we translate into 50 AlphaGeometry problems (we call this set IMO-AG-50). Some problems are split into two due to specifics of our formalization. Figure 8 demonstrates our main result: AlphaGeometry2 solves 42 out of 50 of all 2000-2024 IMO geometry problems, thus surpassing an average gold medallist for the first time [^5] Sinha et al. (2024) previously claimed to achieve a gold medalist performance, but it was done on a subset of IMO problems.. More details are presented in Table 4, which compares various AG2 configurations with other systems, such as AG1 Trinh et al. (2024) and TongGeometry Zhang et al. (2024). We also perform an additional evaluation on a new set of 30 hardest IMO shortlist problems, which are formalizable in the AG2 language, and which have never appeared at IMO. For these additional results see Appendix D.
@@
Wu with AG1 DDAR Sinha et al. (2024) - 21
@@
Wu + AG1 Sinha et al. (2024) - 27
TongGeometry w/o value Zhang et al. (2024) - 28
AG2 with AG1 setup 38 28
TongGeometry full setting Zhang et al. (2024) - 30
AG2 full setting 42 30
Table 4: Evaluation on IMO-AG-50 benchmark. IMO-AG-50 contains all IMO 2000-2024 geometry problems, while IMO-AG-30 introduced in Trinh et al. (2024) contains only a subset formalizable in terms of the AG1 language.
On Figure 7 we present the IMO solve rate as a function of training time (seen tokens during training) for one language model coupled with DDAR via the "classical" tree search described in Section 6. Interestingly, AlphaGeometry2 can already solve 27 out of 50 problems after only 250 training steps with batch size 256, or around 200 million tokens [^6] Note that even without the language model, AlphaGeometry2 can solve 16 problems with its symbolic engine alone (see Figure 8).. We also run ablation studies on how inference settings affect the overall performance (see Figure 9). For a single search tree we find that the optimal configuration is beam size of 128, beam depth of 4 and 32 samples. More samples or a larger beam search does not help solve more problems.
@@
Our geometry experts and IMO medalists consider many AlphaGeometry solutions to exhibit superhuman creativity. In Appendix C we provide several such examples with the detailed analysis. Out of the unsolved IMO problems, we have 2 attempted but not solved and 6 unformalizable problems. The unformalizable problems involve inequalities and variable number of points, which are currently not covered by the AlphaGeometry2 language. Two of the remaining unsolved IMO problems (IMO 2018 P6, IMO 2023 P6) involve advanced geometry problem solving techniques such as inversion, projective geometry or radical axis, which are not implemented in our current DDAR. While such problems in theory can be solved without these techniques, such solutions would require longer inference time, longer proofs and more auxiliary constructions to make up for the lack of the aforementioned machinery in our current DDAR, which hinders AlphaGeometry’s current problem solving capabilities.
@@ 9 Conclusions and Future work
We also presented several studies related to language modeling. First, we showed that our models are quite capable in generating not only auxiliary constructions, but also full proofs, demonstrating a potential for modern language models to operate without external tools, such as symbolic engines. Secondly, we discovered that for AlphaGeometry neither a tokenizer, nor a domain language used to train the model, plays a decisive role. We obtained similar results for custom tokenizers with small vocabularies and the generic large Gemini tokenizer. Training in the domain specific language got us similar results compared to training in natural language. Thirdly, we compared training from scratch and fine-tuning language models pre-trained on math datasets. We found that despite training on the same AlphaGeometry dataset, these models learn slightly different skills, and combining them into our novel search algorithm, Shared Knowledge Ensemble of Search Trees, improves the overall solve rate.
@@
Despite achieving an impressive 84% solve rate on all 2000-2024 IMO geometry problems, there is still a room for improvements. First, our domain language does not allow talking about variable number of points, non-linear equations and problems involving inequalities, which must be addressed in order to fully “solve geometry”. Second, AG2 has not solved all IMO and IMOSL problems. We hypothesize that breaking problems into subproblems and applying Reinforcement learning approaches could close this gap. Finally, in this paper we reported progress on building a fully automated geometry problem solving system, which takes input in natural language and outputs a solution reliably without any hallucinations. Despite good initial results, we think the auto-formalization can be further improved with more formalization examples and supervised fine-tuning.
Acknowledgement
@@
Special thanks to Dawsen Hwang, Edward Lockhart, and Steven Creech for contributions to the development of AlphaGeometry2. We would also like to thank Yifeng Lu, Henryk Michalewski, Ed Chi, David Silver, Pushmeet Kohli, and Demis Hassabis for their thoughtful discussions, help and support.
@@
Appendix A Fine-tuning of math specialized language models on AG data
@@
Even though during the initial transition to AG2, we maintained the AG1 training setup (training from scratch using a custom tokenizer in the AG domain specific language), it is natural to ask whether fine-tuning language models that already possess problem solving capabilities can improve the performance. Such fine-tuning is not immediately possible due to the difference in the utilized tokenizers and training language. In this section we explore the role of the custom tokenizer and the domain specific language, followed by a discussion about fine-tuning of a math specialized Gemini model on AG data.
@@
Tokenizer is an essential part of modern language models and, more broadly, any foundation models [^7] Tokenizer-free models is an active area of research, for example, see Deiseroth et al. (2024). It is generally believed that a tokenizer might be the major bottleneck in the models abilities to do math, e.g. see Singh and Strouse (2024). We investigate this hypothesis in the controlled setting of AlphaGeometry. To do so we train models of the same architecture with different tokenizers: custom tokenizers with vocabularies of a few thousand tokens and the large language model tokenizers with a vocabulary of 300k tokens. Recall that our custom tokenizers are created at word-level, i.e. each token has full meaning, as opposed to subword level tokens. In AG language there are the following types of tokens:
@@
Somewhat surprisingly, we find that AlphaGeometry performance on the 2000-2024 IMO geometry problems stays the same with different tokenizers, which suggests that modern LLM tokenizers might be flexible enough to perform mathematical manipulations.
Domain specific language.
Alongside tokenizers, it is interesting to study the role of the domain specific language in the LM’s ability to solve math problems. It is natural to assume that using domain specific languages simplifies mathematical manipulations and prevents obvious mistakes, which might occur from using a less strict language. To investigate this, we translate all AlphaGeometry2 data from the AlphaGeometry language into natural language and train a new model. Then we compare its performance against the model of the same size trained on the original AlphaGeometry data. Somewhat surprising again, we get the same results on 2000-2024 IMO geometry problems, which opens a path for fine-tuning large language models pre-trained in natural language on math data. Below we demonstrate an example of translating AlphaGeometry into natural language.
@@
Having shown that the custom tokenizer and the domain specific language does not play a critical role for AlphaGeometry, we leverage language models pre-trained on various math data. We start with a Gemini model with 3.3B parameters trained on public math datasets (see Section 7 in Team Gemini (2024)), and fine-tune it in the unsupervised manner on the AlphaGeometry data. On our IMO-AG-50 evaluation set the fine-tuned model performs on par with smaller models and the 3.3B model trained from scratch [^8] We also train even larger models in the supervised manner and achieve the same results. On the other hand, we find that even though all these models are trained on the same AG data, they do produce slightly different auxiliary points proposals and do help each other via the knowledge sharing mechanism described in Section 6, thus forming an ensemble-like system (see Figure 4).
@@
Appendix B Multi-modal
Until now we talked about AG2 as a system that couples a language model together with a symbolic engine. However, since our language model is based on Gemini 1.5, which is multi-modal by design (see Team Gemini (2024)), it is natural to enhance AG model through multi-modal reasoning. For this, we train a new family of models that, alongside the problem text, take the corresponding diagram image as the input. For training and during test time, diagrams are built as described in Section 3.
Despite promising results during the training, we do not observe any improvements in the solve rate on the downstream IMO problems when using this model alone. However, just like in case of fine-tuning pre-trained models (see Section A), we find that the multimodal model produces slightly different auxiliary point proposals. Combined with other models via the knowledge sharing mechanism (see Section 6) this boosts the overall performance. We hypothesize that adding the image on its own might not help that much due to very complicated diagrams, which become very crowded for the IMO problems. The image tokenization process might also play a negative role, as it splits the diagram into independent sequential patches which leads to losing some spatial information. Recall also that some details about the diagram are already provided through the text, as mentioned in Section 7.2, and our symbolic engine, DDAR, does have access to topological aspects of the diagram, e.g. through inspecting sameclock predicates. Furthermore, Chae et al. (2024) shows that vision-language models have poor atomic visual skills, which suggests that adding visual elements might not aid the geometry problem-solving process. Finally note that the core of geometry problem-solving lies in algebraic reasoning, as previously demonstrated in Trinh et al. (2024), rather than geometric reasoning. Many human IMO contestants can reliably solve geometry problems (including very hard problems such as IMO 2011 P6) using computational methods such as complex numbers, barycentric coordinates and trigonometry bashing, which means that visual information and diagrams are not critical to solving geometry problems.
Appendix C Featured AlphaGeometry2 solutions

### v3

@@ 6.2 Inference setup
A new problem is solved via the search algorithm described in section 5 with multiple search trees and multiple language models of different sizes. In contrast to AG1, we use top-k sampling with temperature $t=1.0$ and $k=32$. Note that a high temperature and multiple samples are essential for solving IMO problems. With the greedy decoding $t=0.0,k=1$, and no tree search, our models can solve only two problems out of 26 that require auxiliary constructions. Increasing the temperature to $t=1.0$ and using $k=32$ samples (without a search tree) allows our language models to solve 9 out of 26 problems. Lower temperatures $t<1.0$ do not produce diverse enough auxiliary constructions (see Figure 7), while higher temperatures result in an increasing number LM outputs with a wrong domain language syntax.
@@
Note that by definition, $S_{1}\subset S_{2}\subset S_{3}$. Once these three sets are computed, we serialize and concatenate them into a string called analysis string, using our domain-specific language. This string is fed into the LM, together with the original problem statement as follows: <problem_statement> serialized($S_{1}$) serialized($S_{2}-S_{1}$) serialized($S_{3}-S_{2}$). In contrast, the input to the AG1 LM is simply <problem_statement>.
@@ 7 Results
Our main downstream metric is the solve rate on IMO geometry problems. There are a total of 45 geometry problems in 2000-2024 IMO, which we translate into 50 AlphaGeometry problems (we call this set IMO-AG-50). Some problems are split into two due to the specifics of our formalization. Figure 8 demonstrates our main result: AlphaGeometry2 solves 42 out of 50 of all 2000-2024 IMO geometry problems, thus surpassing an average gold medallist for the first time555sinha2024wusmethodboostsymbolic previously claimed to achieve a gold medalist performance, but it was done on a subset of IMO problems.. More details are presented in Table 4, which compares various AG2 configurations with other systems, such as AG1 Trinh2024 and TongGeometry zhang2024proposing. We also perform an additional evaluation on a new set of 30 hardest IMO shortlist problems, which are formalizable in the AG2 language, and which have never appeared at IMO. For these additional results, see Appendix E.
@@
Wu with AG1 DDAR sinha2024wusmethodboostsymbolic - 21
@@
Wu + AG1 sinha2024wusmethodboostsymbolic - 27
TongGeometry w/o value zhang2024proposing - 28
AG2 with AG1 setup (a single search tree) 38 28
TongGeometry full setting zhang2024proposing - 30
AG2 full setting (multiple search trees) 42 30
Table 4: Evaluation on IMO-AG-50 benchmark. IMO-AG-50 contains all IMO 2000-2024 geometry problems, while IMO-AG-30 introduced in Trinh2024 contains only a subset formalizable in terms of the AG1 language.
On Figure 7, we present the IMO solve rate as a function of training time (seen tokens during training) for one language model coupled with DDAR via the "classical" tree search described in Section 5. Interestingly, AlphaGeometry2 can already solve 27 out of 50 problems after only 250 training steps with a batch size of 256, or around 200 million tokens [^6] Note that even without the language model, AlphaGeometry2 can solve 16 problems with its symbolic engine alone (see Figure 8).. We also run ablation studies on how inference settings affect the overall performance (see Figure 9). For a single search tree, we find that the optimal configuration is a beam size of 128, a beam depth of 4, and 32 samples. More samples or a larger beam search do not help solve more problems.
@@
Our geometry experts and IMO medalists consider many AlphaGeometry solutions to exhibit superhuman creativity. In Appendix D we provide several such examples with the detailed analysis. Out of the unsolved IMO problems, we have 2 attempted but not solved and 6 unformalizable problems. The unformalizable problems involve inequalities and variable number of points, which are currently not covered by the AlphaGeometry2 language. Two of the remaining unsolved IMO problems (IMO 2018 P6, IMO 2023 P6) involve advanced geometry problem-solving techniques such as inversion, projective geometry or radical axis, which are not implemented in our current DDAR. While such problems in theory can be solved without these techniques, such solutions would require longer inference time, longer proofs and more auxiliary constructions to make up for the lack of the aforementioned machinery in our current DDAR, which hinders AlphaGeometry’s current problem-solving capabilities.
@@ 8 Conclusions and Future Work
We also present several studies related to language models in general. First, in Appendix B we demonstrate that for AlphaGeometry neither a tokenizer, nor a domain language used to train the model, plays a decisive role. We obtained similar results for custom tokenizers with small vocabularies and the generic large Gemini tokenizer. Training in the domain-specific language achieves similar results compared to training in natural language. Second, in the same appendix, we compare training from scratch and fine-tuning language models pre-trained on math datasets. We find that despite training on the same AlphaGeometry dataset, these models learn slightly different skills, and combining them into our novel search algorithm, Shared Knowledge Ensemble of Search Trees, improves the overall solve rate. Third, in Appendix F, we show that our models are quite capable in generating not only auxiliary constructions, but also full proofs, demonstrating a potential for modern language models to operate without external tools, such as symbolic engines. Finally, in Appendix G we show that AlphaGeometry2 can be used to build a fully automated geometry problem-solving system that takes inputs in natural language and reliably outputs corresponding diagrams and full solutions without any hallucinations.
@@
Despite achieving an impressive 84% solve rate on all 2000-2024 IMO geometry problems, there is still room for improvement. First, our domain language does not allow talking about variable number of points, non-linear equations, and problems involving inequalities, which must be addressed in order to fully “solve the Euclidean geometry”. Extending AlphaGeometry to encompass these topics is a substantial undertaking that falls beyond the scope of this work. To illustrate the complexity involved, in Appendix H we present the list of inequality rules the symbolic engine would need to incorporate. Second, AG2 has not solved all IMO and IMOSL problems. We hypothesize that breaking problems into subproblems and applying Reinforcement learning approaches could close this gap.
Code availability
Code for the Python implementation of the symbolic engine (DDAR2), along with multiple examples of proven IMO problems, will be shared at https://github.com/google-deepmind/alphageometry2. Code for training Gemini models is not provided because it uses internal Google infrastructure. Code for running the full tree search with multiple Gemini models is not provided because it relies on internal Google infrastructure and running many workers in parallel.
Data availability
We share 27 IMO problems translated to the AlphaGeometry language along with their diagrams and solutions. They can be found in the file test.py within the provided repository.
Acknowledgments
@@
Special thanks to Dawsen Hwang, Edward Lockhart, and Steven Creech for contributions to the development of AlphaGeometry2. We would also like to thank Yifeng Lu, Henryk Michalewski, Ed Chi, David Silver, Pushmeet Kohli, and Demis Hassabis for their thoughtful discussions and support.
Given a random diagram, if DDAR proves: Then let X = Then if X is nonempty, we create a synthetic proof that says: "when X moves, … Auxiliary constructions will be the following points and everything else they depend on Case
cong a b c d P(b,c,d)P(a)P(b,c,d)-P(a) circle center $b$, radius $cd$ goes through a fixed point $a$ 2
P(a)P(b,c,d)P(a)-P(b,c,d) $a$ moves on a fixed circle $b,c,d$ 8
P(a,b)P(c,d)P(a,b)-P(c,d) the distance between $a$ and $b$ is fixed $c,d$ 9
cong a b a c P(a)P(b,c)P(a)-P(b,c) $a$ moves on a fixed line $b,c$ 7
$P(b,c)-P(a)\geq M$ $b$ & $c$ are equidistant to a fixed point, when M move $a$ 4
cyclic a b c d P(b,c,d)P(a)P(b,c,d)-P(a) the circumcircle of $b$ $c$ $d$ moves through a fixed point $a$ 1
P(d)P(a,b,c)P(d)-P(a,b,c) $d$ moves on a fixed circle $a,b,c$ 8
coll a b c P(b,c)P(a)P(b,c)-P(a) line $b$ $c$ goes through a fixed point $a$ 3
P(c)P(a,b)P(c)-P(a,b) $c$ moves on a fixed line $a,b$ 7
eqangle b a b c e d e f P(a,b,c)P(d,e,f)P(a,b,c)-P(d,e,f) the angle $a$ $b$ $c$ has a fixed value $d,e,f$ 11
P(b)P(a,c,d,e,f)P(b)-P(a,c,d,e,f) the point $b$ moves on a fixed circle $a,c,d,e,f$ 8
para a b c d P(b,c,d)P(a)P(b,c,d)-P(a) The line through $b$ and $\parallel cd$ moves through a fixed point $a$ 5
P(c,d)P(a,b)P(c,d)-P(a,b) The line $cd$ is always parallel to a fixed line $a,b$ 10
P(a)P(b,c,d)P(a)-P(b,c,d) $a$ moves on a fixed line $b,c,d$ 7
perp a b c d P(b,c,d)P(a)P(b,c,d)-P(a) the line through $b$ and $\perp cd$ moves through fixed point $a$ 6
P(c,d)P(a,b)P(c,d)-P(a,b) the line $cd$ is always $\perp$ to a fixed line $a,b$ 10
P(a)P(b,c,d)P(a)-P(b,c,d) $a$ moves on a fixed line $b,c,d$ 7
Table 5: 17 cases where locus-type statements are detected during data generation. These 17 cases produce 11 different types of locus-type statements, as numbered in the last column.
Appendix A Related work
Neuro-symbolic system: At the high level, AlphaGeometry is a neuro-symbolic system with a generative component (i.e. language model). This framework was motivated from the fact that a major limitation of automated theorem provers is the ability to generate original mathematical terms, which could be addressed using language models polu2020generativelanguagemodelingautomated. Such framework was used in several papers such as Trinh2024; wei2024provingineq; kaiyu2025inequalities to great success. Additionally, because the symbolic system is able to provide consistent and verifiable feedback, neuro-symbolic systems have been augmented with reinforcement learning techniques to give AI agents the ability to self-improve without human intervention jha2024rlsfreinforcementlearningsymbolic; peng-etal-2023-geodrl. In contrast, our AG2 does not use any reinforcement learning techniques to achieve gold-medalist level performance on IMO geometry problems.
LLM-based feedback: In the era of large language models (LLMs), one might think of using them (instead of symbolic engines) as verifiers due to their versatility. This approach is used in shinn2023reflexionverbalrl; madaan2023selfrefine; chen2024teachingllmselfdebug; lightman2024letsverify; zhang2025generativeverifiers. However, this approach lies on the assumption that LLMs are reliable verifiers and this notion has been challenged huang2024llmcannotselfcorrect; stechly2025onselfverificationlimitation. Solving Olympiad geometry problems involves performing numerous precise algebraic manipulations consistently throughout a potentially long solution, whereas LLMs are notoriously unreliable even for basic arithmetics yan2025phdlevelllmstrulygrasparithmetic. And so, it is natural for AlphaGeometry to favor symbolic engines over LLMs to provide consistent feedback to the solver.
Visual reasoning: While AlphaGeometry can solve difficult Olympiad geometry problems and surpass the performance of top human experts, many papers have demonstrated the limitations of various foundation models in performing geometric reasoning mouselinos-etal-2024-geometric-reasoning-gap; wang2025dollmunderstandgeometry. As visual thinkers Tversky2009-think-with-sketch, humans want to teach AI agents to leverage visual information to further unlock its reasoning capabilities hu2024visualsketchpad. Even though AlphaGeometry mostly relies on algebraic reasoning, we discuss incorporating visual modality in Appendix C.
@@ Appendix A Related work
Appendix B Fine-tuning of math specialized language models on AG data
@@
Even though during the initial transition to AG2, we maintained the AG1 training setup (training from scratch using a custom tokenizer in the AG domain-specific language), it is natural to ask whether fine-tuning language models that already possess problem-solving capabilities can improve the performance. Such fine-tuning is not immediately possible due to the difference in the utilized tokenizers and training language. In this section, we explore the role of the custom tokenizer and the domain-specific language, followed by a discussion about fine-tuning of a math-specialized Gemini model on AG data.
@@
Tokenizer is an essential part of modern language models and, more broadly, any foundation models [^7] Tokenizer-free models is an active area of research, for example, see deiseroth2024tfreetokenizerfreegenerativellms. It is generally believed that a tokenizer might be the major bottleneck in the models’ abilities to do math, e.g. see singh2024tokenizationcountsimpacttokenization. We investigate this hypothesis in the controlled setting of AlphaGeometry. To do so, we train models of the same architecture with different tokenizers: custom tokenizers with vocabularies of a few thousand tokens and the large language model tokenizers with a vocabulary of 300k tokens. Recall that our custom tokenizers are created at word-level, i.e. each token has full meaning, as opposed to subword-level tokens. In AG language, there are the following types of tokens:
@@
Somewhat surprisingly, we find that AlphaGeometry’s performance on the 2000-2024 IMO geometry problems stays the same with different tokenizers, which suggests that modern LLM tokenizers might be flexible enough to perform mathematical manipulations.
domain-specific language.
Alongside tokenizers, it is interesting to study the role of the domain-specific language in the LM’s ability to solve math problems. It is natural to assume that using domain-specific languages simplifies mathematical manipulations and prevents obvious mistakes, which might occur from using a less strict language. To investigate this, we translate all AlphaGeometry2 data from the AlphaGeometry language into natural language and train a new model. Then we compare its performance against the model of the same size trained on the original AlphaGeometry data. Somewhat surprising again, we get the same results on 2000-2024 IMO geometry problems, which opens a path for fine-tuning large language models pre-trained in natural language on math data. Below, we demonstrate an example of translating AlphaGeometry into natural language.
@@
Having shown that the custom tokenizer and the domain-specific language does not play a critical role for AlphaGeometry, we leverage language models pre-trained on various math data. We start with a Gemini model with 3.3B parameters trained on public math datasets (see Section 7 in team2024gemini), and fine-tune it in an unsupervised manner on the AlphaGeometry data. On our IMO-AG-50 evaluation set, the fine-tuned model performs on par with smaller models and the 3.3B model trained from scratch [^8] We also train even larger models in a supervised manner and achieve the same results. On the other hand, we find that even though all these models are trained on the same AG data, they do produce slightly different auxiliary points proposals and do help each other via the knowledge sharing mechanism described in Section 5, thus forming an ensemble-like system (see Figure 4).
@@
Appendix C Multi-modal
Until now, we have talked about AG2 as a system that couples a language model together with a symbolic engine. However, since our language model is based on Gemini 1.5, which is multi-modal by design (see team2024gemini), it is natural to enhance the AG system through multi-modal reasoning. For this, we train a new family of models that, alongside the problem text, take the corresponding diagram image as the input.
Despite promising results during the training, we do not observe any improvements in the solve rate on the downstream IMO problems when using this model alone. However, just like in the case of fine-tuning pre-trained models (see Section B), we find that the multimodal model produces slightly different auxiliary point proposals. Combined with other models via the knowledge sharing mechanism (see Section 5) this boosts the overall performance. We hypothesize that adding the image on its own might not help that much due to very complicated diagrams, which become very crowded for the IMO problems. The image tokenization process might also play a negative role, as it splits the diagram into independent sequential patches, which leads to the loss of some spatial information. Recall also that some details about the diagram are already provided through the text, as mentioned in Section 6.2, and our symbolic engine, DDAR, does have access to topological aspects of the diagram, e.g., through inspecting sameclock predicates. Furthermore, chae2024decomposingvlm shows that vision-language models have poor atomic visual skills, which suggests that adding visual elements might not aid the geometry problem-solving process. Finally, note that the core of geometry problem-solving lies in algebraic reasoning, as previously demonstrated in Trinh2024, rather than geometric reasoning. Many human IMO contestants can reliably solve geometry problems (including very hard problems such as IMO 2011 P6) using computational methods such as complex numbers, barycentric coordinates and trigonometry bashing, which means that visual information and diagrams are not critical to solving geometry problems.
Appendix D Featured AlphaGeometry2 solutions